Caleb is 54 years younger than Rebecca. 7 years ago, Rebecca's age was 4 times Caleb's age. How old is Caleb now?
Answer by stanbon(75887)

Caleb is 54 years younger than Rebecca. 7 years ago, Rebecca's age was 4 times Caleb's age. How old is Caleb now?
--------------
Equations:
c = r-54
r-7 = 4(c-7)
--------------------
Note: r = c+54
----------------------------
Substitute for "r" and solve for "c":
c+54-7 = 4c-28
c+47 = 4c-28
3c = 75
Caleb = 25 yrs (Caleb's age now)
===========
Cheers,
Stan H.
===========